A bag contains 2 red, 3 green and 2 blue balls. Two balls are drawn at random. What is the probability that none of the balls drawn is blue? 


1
Expert's answer
2021-06-14T15:25:04-0400

The total number of balls is


"2+3+2=7"


The number of balls which are not blue is "2+3=5."

The probability that the first ball is not blue and the second ball is bot blue is


"P(B'B')=\dfrac{5}{7}(\dfrac{5-1}{7-1})=\dfrac{10}{21}"

The probability that none of the balls drawn is blue is "\dfrac{10}{21}" .